Make New Entry,
Make Followup Entry
User name Chiranjib
Log entry time 07:26:13 on November 2,2008
Entry number 246193
keyword=NMR at the end of the shift
Beam is still not back yet. I tried to do an NMR in the optics position for the pumping chamber. The VI got stuck and I called Jin . He remotely did that while I rebooted the wavetek 80 trigger.
the NMR signal in the pumping chamber read 38 mV roughly which corresponds to a polarization number of